# Graph Flow Matching: Enhancing Image Generation with Neighbor-Aware Flow Fields

### v2

@@ 5 Conclusion
We introduced Graph Flow Matching (GFM), a lightweight architectural enhancement that improves the expressiveness of flow matching generative models by incorporating local neighborhood structure via a graph-based “diffusion” term. Through a reaction–diffusion decomposition, GFM enables pointwise velocity predictors to aggregate contextual information from neighboring samples along the flow trajectory—offering a principled and scalable mechanism for improving the interpolation of the flow field.
@@ Appendix A Appendix
This appendix provides supplementary material including implementation details, the theoretical justification for Lemma 3.1 and extended experimental results. We organize it into the following sections:
  • Section B: Proof of Lemma 3.1
  • Section C: Experimental details
  • Section D: Diffusion-reaction with KNN ablation study
  • Section E: Image generation timing analysis
  • Section F: Comparison against leading generative models.
  • Section G: Additional generated samples
Appendix B Proof of Lemma 3.1
Proof.
The inequality follows from the function space representable by $v_{\theta}=v_{\text{react}}+v_{\text{diff}}$ strictly including the space representable by $v_{\text{react}}$ alone [50]. At minimum, $v_{\text{diff}}$ could learn to be zero, recovering $v_{\text{react}}$’s performance. For data with local coherence, where velocity varies smoothly across the manifold (common in natural images [51]), graph neural networks effectively exploit this structure [52, 53]. The graph diffusion term leverages local structure through message passing, allowing each point to aggregate information from its neighbors to better capture the local geometry of the flow. ∎
This completes the proof that the inequality is strict when the data exhibits local coherence.
Appendix C Experimental Details
@@
We implemented our code using PyTorch (offered under BSD-3 Clause license) and conducted experiments on Nvidia hardware (RTX 4090 for ADM UNet variants and RTX A6000 for DiT-L/2 experiments). All models from [6] were evaluated under identical conditions, with experiment tracking via Weights and Biases (wandb). The following subsections detail our datasets, architectures, and hyperparameter settings.
C.1 Datasets
@@
We follow the exact preprocessing pipelines of [6] for all datasets, including VAE encoding at $256\times 256$ resolution. To characterize the intrinsic variability of each dataset, we compute an internal FID score by randomly splitting each dataset into two non-overlapping halves and measuring the FID between them. Higher internal FID values indicate greater intrinsic variability within the dataset (Table 4).
@@
The LSUN Church dataset [8] comprises approximately 126,227 training images of outdoor church buildings and architectural structures. This dataset presents particular challenges due to the complexity of architectural details and structural coherence required in the generated images.
@@
The LSUN Bedroom dataset [8] contains around 3,033,042 training images of indoor bedroom scenes, making it substantially larger than the other datasets used in our experiments. This dataset features complex indoor environments with various furniture arrangements, lighting conditions, and textures, requiring models to capture both local details and global room layouts.
@@
The Flickr-Faces-HQ (FFHQ) dataset [9] consists of 70,000 high-quality human portrait photographs. This dataset is particularly challenging for generative models as human faces contain subtle details that are easily perceived when incorrectly generated, making it an excellent benchmark for evaluating generative fidelity.
@@
The AFHQ-Cat subset [10] contains $\sim$5,000 training images of cat faces from the Animal Faces-HQ dataset. Despite its smaller size compared to the other datasets, it presents unique challenges in capturing fine details of animal features and fur textures.
@@
The CelebA-HQ dataset [11] provides 30,000 celebrity face images. This dataset is a high-quality version of the original CelebA dataset, with improved resolution and reduced compression artifacts. CelebA-HQ features greater diversity in facial attributes, expressions, and backgrounds compared to FFHQ, though with a stronger bias toward frontal-facing poses and celebrity appearances.
@@
C.2 Training Algorithm
We train our latent flow matching models as in LFM ([6]), where the authors used the convention that noise is at $t=1$ and the target distribution is at $t=0$, which we summarize here. At each iteration, a real image $x_{0}$ from the data distribution $p_{0}$ is sampled and encoded into latent space via the VAE encoder $E$, yielding $z_{0}=E(x_{0})$. A sample from a standard Gaussian latent $z_{1}\sim\mathcal{N}(0,I)$ is then taken and a single time $t\sim\mathcal{U}(0,1)$ is drawn. The interpolation point $z_{t}=(1-t)z_{0}+tz_{1}$ lies along the straight-line transport plan between $z_{0}$ and $z_{1}$. Then, the instantaneous velocity $v_{\theta}(z_{t},t)$ is predicted with the flow network and the squared-error loss against the true transport increment $z_{1}-z_{0}$ is computed. Finally, all network parameters are updated by gradient descent on this loss.
C.3 Network configurations
The network configurations for $v_{\text{react}}$ for ADM from [6] are provided in Table 6 and that for DiT-L/2 can be found in [6] and in the corresponding GitHub repository. [^2] https://github.com/VinAIResearch/LFM/tree/main Those for $v_{\text{diff}}$ are provided in Footnote 4 and Table 7.
C.3.1 MPNN Diffusion Architecture
The MPNN diffusion module implements the velocity correction term $v_{\text{diff}}$ using a message-passing neural network architecture inspired by the connection between graph neural networks and partial differential equations. As described in Equation 8, our MPNN architecture models the diffusion term as:
$v_{\text{diff}}(\mathbf{x}_{t},t)=-N_{\theta}^{(2)}\left(\mathbf{G}^{\top}\left[\sigma\left(\mathbf{G}\,N_{\theta}^{(1)}(\mathbf{x}_{t},t)\right)\right],\,t\right),$
where $\mathbf{G}$ is the graph gradient operator that computes differences between connected node features, $\sigma$ is a nonlinearity (we used the ELU function in all our experiments), and $N_{\theta}^{(1)}$ and $N_{\theta}^{(2)}$ are identical U-Net architectures (from [44], with code obtained from [54]) whose architectural hyperparameters are shown in Table 4. The graph gradient $\mathbf{G}$ is constructed dynamically using attention-based adjacency, enabling the model to adapt its neighborhood structure based on the input latent representations.
Table 5: Architectural hyperparameters for models using MPNN as the diffusion term ($v_{\text{diff}}$). Both $N_{\theta}^{(1)}$ and $N_{\theta}^{(2)}$ use identical U-Net architectures from [44] [^4] Implementation adapted from https://github.com/annegnx/PnP-Flow with the parameters shown.
Dataset $v_{\text{react}}$ Base Channels Channel Mult Res Blocks Attn. Resolutions
@@ C.3.1 MPNN Diffusion Architecture
C.3.2 GPS Diffusion Architecture
The GPS diffusion module implements the velocity correction term $v_{\text{diff}}$ using a graph transformer architecture as described in Section 3. We adapted the implementation from the PyTorch Geometric repository [^5] https://github.com/pyg-team/pytorch_geometric/blob/master/examples/graph_gps.py. Table 7 presents the hyperparameters used for each dataset in our GPS module experiments. The module:
@@ C.3.2 GPS Diffusion Architecture
Table 6: ADM $v_{\text{react}}$ configurations taken from [6]. The CelebA-HQ configuration is the one obtained by testing the corresponding checkpoint from [6].
@@
Table 7: Architectural hyperparameters for GPS-based diffusion term ($v_{\text{diff}}$). All models used multihead attention with $4$ heads and time embeddings with $8$ learnable frequencies.
Dataset $v_{\text{react}}$ Hidden Dim Layers PE Dim Walk Length Time Channels
@@
C.4 Training Hyperparameters
We evaluated GFM using two reaction term architectures: DiT-L/2 and ADM U-Net. For both, we initialized $v_{\text{react}}$ with corresponding architectures and checkpoints from [6], then jointly trained the reaction and graph correction (“diffusion term”) components. For AFHQ-Cat experiments, we used the CelebA-HQ checkpoints to initialize the reaction terms. For evaluation of performance metrics, FID scores were monitored every $3$K iterations, with training terminating after $80$K iterations without improvement. We saved models with the best FID scores and evaluated them using $50$K generated samples. For CelebA-HQ comparisons, we obtained higher, i.e., worse FID scores than reported in [6] when evaluating their provided checkpoints on our machines [^6] https://github.com/VinAIResearch/LFM/tree/main. To establish fair baselines, we trained these checkpoints for an additional $100$K iterations using the training settings in [6] and used the models with the lowest FID scores as the baseline models for the corresponding DiT and ADM experiments.
Tables 8 and 9 detail the training hyperparameters for our GFM variants with DiT-L/2 and ADM U-Net backbones. All models used AdamW optimization with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The original training settings for the [6] checkpoints are provided in Tables 10 and 11. Similar to [6], a cosine annealing scheduler was used for the learning rate. For ablation studies in Table 3 with identity adjacency matrices, all other settings (besides their adjacency matrices) remained identical to their corresponding graph-based experiments.
Table 8: Training hyperparameters when $v_{\text{react}}$ is DiT-L/2 across datasets. All models used AdamW optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$.
$v_{\text{diff}}$ Parameter LSUN Church FFHQ LSUN Bedroom AFHQ-Cat CelebA-HQ
@@ C.4 Training Hyperparameters
Batch size $42$ $32$ $32$ $20$ $32$
Training iter. $40$K $40$K $60$K $60$K $60$K
# GPUs $1$ $1$ $1$ $1$ $1$
@@
Training iter. $40$K $40$K $40$K $60$K $60$K
# GPUs $1$ $1$ $1$ $1$ $1$
Table 9: Training hyperparameters when $v_{\text{react}}$ is ADM [6] across datasets. All models used AdamW optimizer with $\beta_{1}=0.9,\beta_{2}=0.999$.
$v_{\text{diff}}$ Parameter LSUN Church FFHQ LSUN Bedroom AFHQ-Cat CelebA-HQ
@@
Training iter. $50$K $120$K $40$K $200$K $60$K
# GPUs $1$ $1$ $1$ $1$ $1$
@@
Training iter. $50$K $120$K $40$K $200$K $60$K
# GPUs $1$ $1$ $1$ $1$ $1$
Table 10: Training hyperparameters for the ADM checkpoints used as $v_{\text{react}}$, reproduced exactly from [6].
@@
# GPUs 2 1 1
# training days 1.3 4.4 6.9 & 7.3
Table 11: Training hyperparameters for the DiT checkpoints used as $v_{\text{react}}$, reproduced exactly from [6].
@@
# GPUs 1 1 2 & 1
# training days 5.8 5.5 6.1 & 12.5
Appendix D Ablation Study: Diffusion-Reaction with KNN
In this ablation, we swap out our attention-based adjacency for a simple $k$-nearest-neighbour (KNN) graph in latent space: for each node, we connect to its $k$ closest latents and assign each edge a learnable weight, so that in the diffusion term, information is aggregated strictly from a node’s top-$k$ neighbors instead of via attention scores. Concretely, we implement the diffusion term as a graph Laplacian multiplied by a per-channel diffusivity vector, mirroring the physical reaction–diffusion system in Equation 6. The overall velocity field remains the sum of the original flow-matching “reaction” network and this KNN-based “diffusion” operator. Let $X\in\mathbb{R}^{B\times C\times H\times W}$ be the batch of latent tensors at time $t$. This network implements:
$\frac{dX}{dt}(t)\;=\;v_{\mathrm{react}}\bigl{(}X(t),\,t\bigr{)}\;+\;\underbrace{\kappa_{\theta}\bigl{(}X(t),\,t\bigr{)}}_{\begin{subarray}{c}\text{per-channel}\\ \text{diffusivity}\end{subarray}}\;\odot\;\Bigl{[}L(X(t))\,X_{\mathrm{flat}}(t)\Bigr{]}\,,$ (10)

### v3

@@ 5 Conclusion
We introduced Graph Flow Matching (GFM), a lightweight architectural enhancement that improves the expressiveness of flow matching generative models by incorporating local neighborhood structure via a graph-based “diffusion” term. Through a reaction–diffusion decomposition, GFM enables pointwise velocity predictors to aggregate contextual information from neighboring samples along the flow trajectory, offering a principled and scalable mechanism for improving the interpolation of the flow field.
@@ Appendix A Appendix
This appendix provides supplementary material including further discussion, implementation details, and extended experimental results. We organize it into the following sections:
  • Section A.1: Further discussion on the motivation for our framework.
  • Section B: Experimental details
  • Section C: Diffusion-reaction with KNN ablation study
  • Section D: Image generation timing analysis
  • Section E: Comparison against leading generative models.
  • Section F: Additional generated samples
A.1 From Interpolation to Neighbor‑Aware Velocity Prediction
The smoothness of the flow field $\mathbf{v}(\mathbf{x},t)$ is a well established property in the literature. Under standard regularity conditions, such as a smooth Gaussian base distribution and a target distribution with finite moments, the learned flow field $\mathbf{v}(\mathbf{x},t)$ is typically Lipschitz continuous in both $\mathbf{x}$ and $t$, ensuring bounded gradients and stable interpolation trajectories (gao2024gaussian; gao2024convergence). In theoretical analyses, sufficient regularity in $(\mathbf{x},t)$ is often assumed to establish convergence guarantees (fukumizu2024flow). Additionally, the learned vector field satisfies the continuity equation, ensuring mass-conserving evolution of the probability density over time (lipman2023flowmatchinggenerativemodeling; albergo2022building; gao2024gaussian).
Viewed through this lens, flow matching may be interpreted as a scattered-data interpolation problem: the model observes velocities at a finite set of input points $(\mathbf{x}_{t_{i}},t_{i})$ and must reconstruct a smooth vector field over the entire domain. This perspective parallels classical interpolation methods in numerical analysis, such as radial basis function kernels (micchelli1984interpolation; wendland2004scattered), splines (de1978practicalSplines), or moving least squares (levin1998_MLSapproximation), which all rely on smoothness and local support to stabilize predictions.
In these methods, incorporating local neighbor information, typically via compact support functions or interpolation stencils, improves approximation accuracy and generalization. For instance, radial basis function interpolation computes values at a query point as a weighted sum over nearby samples. Inspired by this, our neighbor-aware velocity module aggregates information from neighboring latent points via a graph-based message-passing mechanism, thereby instantiating a comparable inductive bias through a parameterized, data-driven mechanism.
Contemporary flow matching models often adopt a pointwise prediction scheme, directly regressing from $(\mathbf{x}_{t},t)$ to $\mathbf{v}(\mathbf{x}_{t},t)$ without modeling relationships across neighboring points. This resembles early point cloud models such as PointNet (qi2017pointnet), which lacked mechanisms to capture local geometric structure, a limitation addressed by PointNet++ (qi2017pointnetplusplus) through hierarchical neighborhood aggregation, and by DGCNN (DGCNN_paper), which dynamically constructs graphs to model local relationships via edge convolutions. These architectures demonstrated the benefits of incorporating local context, an idea we adapt for velocity field prediction.
In high-dimensional settings, the absence of neighborhood modeling can hinder generalization, particularly in sparsely supervised, high curvature, or multimodal regions (Zhang_sparseSubspaceLearning; Folco_di2025highDLearning). Given the smoothness of the velocity field $\mathbf{v}(\mathbf{x},t)$, points that are proximate in latent space may lie within the same semantic region and follow similar flow trajectories. Just as nearby points in classification tend to share labels, points close in $\mathcal{X}\times[0,1]$ may exhibit similar velocities. Leveraging this local correlation may help mitigate issues that arise in undersampled or complex regions.
Figure 3: LSUN Bedroom and LSUN Church samples ($256\times 256$) generated using the same random seed by: (top) baseline DiT-L/2 (dao2023flow), (middle) DiT-L/2 with MPNN-based correction, and (bottom) DiT-L/2 with GPS-based graph correction (GPS_rampavsek2022recipe). GFM variants generate more complete spatial structures and sharper boundaries compared to the baseline model.
Appendix B Experimental Details
@@
We implemented our code using PyTorch (offered under BSD-3 Clause license) and conducted experiments on Nvidia hardware (RTX 4090 for ADM UNet variants and RTX A6000 for DiT-L/2 experiments). All models from (dao2023flow) were evaluated under identical conditions, with experiment tracking via Weights and Biases (wandb). The following subsections detail our datasets, architectures, and hyperparameter settings.
B.1 Datasets
@@
We follow the exact preprocessing pipelines of (dao2023flow) for all datasets, including VAE encoding at $256\times 256$ resolution. To characterize the intrinsic variability of each dataset, we compute an internal FID score by randomly splitting each dataset into two non-overlapping halves and measuring the FID between them. Higher internal FID values indicate greater intrinsic variability within the dataset (Table 4).
@@
The LSUN Church dataset (LSUN_dataset_yu15lsun) comprises approximately 126,227 training images of outdoor church buildings and architectural structures. This dataset presents particular challenges due to the complexity of architectural details and structural coherence required in the generated images.
@@
The LSUN Bedroom dataset (LSUN_dataset_yu15lsun) contains around 3,033,042 training images of indoor bedroom scenes, making it substantially larger than the other datasets used in our experiments. This dataset features complex indoor environments with various furniture arrangements, lighting conditions, and textures, requiring models to capture both local details and global room layouts.
@@
The Flickr-Faces-HQ (FFHQ) dataset (FFHQ_dataset_karras2019style) consists of 70,000 high-quality human portrait photographs. This dataset is particularly challenging for generative models as human faces contain subtle details that are easily perceived when incorrectly generated, making it an excellent benchmark for evaluating generative fidelity.
@@
The AFHQ-Cat subset (AFHQ_dataset_choi2020starganv2) contains $\sim$5,000 training images of cat faces from the Animal Faces-HQ dataset. Despite its smaller size compared to the other datasets, it presents unique challenges in capturing fine details of animal features and fur textures.
@@
The CelebA-HQ dataset (CelebAHQ_dataset_karras2018progressive) provides 30,000 celebrity face images. This dataset is a high-quality version of the original CelebA dataset, with improved resolution and reduced compression artifacts. CelebA-HQ features greater diversity in facial attributes, expressions, and backgrounds compared to FFHQ, though with a stronger bias toward frontal-facing poses and celebrity appearances.
@@
B.2 Training Algorithm
We train our latent flow matching models as in LFM ((dao2023flow)), where the authors used the reverse convention that noise is at $t=1$ and the target distribution is at $t=0$, which we summarize here. At each iteration, a real image $\mathbf{i}_{0}$ from the data distribution $p_{0}$ is sampled and encoded into latent space via the VAE encoder $E$, yielding $\mathbf{x}_{0}=E(\mathbf{i}_{0})$. A sample from a standard Gaussian latent $\mathbf{x}_{1}\sim\mathcal{N}(0,I)$ is then taken and a single time $t\sim\mathcal{U}(0,1)$ is drawn. The interpolation point $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}$ which lies along the straight‐line transport plan between $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ is constructed. Then, the instantaneous velocity $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ is predicted with the flow network and the squared‐error loss against the “true” transport velocity $\mathbf{x}_{1}-\mathbf{x}_{0}$ is computed. Finally, all network parameters are updated by gradient descent on this loss.
B.3 Network configurations
The network configurations for $\mathbf{v}_{\text{react}}$ for ADM from (dao2023flow) are provided in Table 6 and that for DiT-L/2 can be found in (dao2023flow) and in the corresponding GitHub repository. [^2] https://github.com/VinAIResearch/LFM/tree/main Those for $\mathbf{v}_{\text{diff}}$ are provided in Table 5 and Table 7.
MPNN Diffusion Architecture
The MPNN diffusion module implements the velocity correction term $\mathbf{v}_{\text{diff}}$ using a message-passing neural network architecture inspired by the connection between graph neural networks and partial differential equations. As described in Equation 8, our MPNN architecture models the diffusion term as:
$\mathbf{v}_{\text{diff}}(\mathbf{x}_{t},t)=-N_{\theta}^{(2)}\left(\mathbf{G}^{\top}\left[\sigma\left(\mathbf{G}\,N_{\theta}^{(1)}(\mathbf{x}_{t},t)\right)\right],\,t\right),$
where $\mathbf{G}$ is the graph gradient operator that computes differences between connected node features, $\sigma$ is a nonlinearity (we used the ELU function in all our experiments), and $N_{\theta}^{(1)}$ and $N_{\theta}^{(2)}$ are identical U-Net architectures (from (PnPUNet_huang2021variational), with code obtained from (pnpFlow_martin2025pnpflow)) whose architectural hyperparameters are shown in Table 5. The graph gradient $\mathbf{G}$ is constructed dynamically using attention-based adjacency, enabling the model to adapt its neighborhood structure based on the input latent representations.
Dataset $\mathbf{v}_{\text{react}}$ Base Channels Channel Mult Res Blocks Attn. Resolutions
@@ B.3 Network configurations
Table 5: Architectural hyperparameters for models using MPNN as the diffusion term ($\mathbf{v}_{\text{diff}}$). Both $N_{\theta}^{(1)}$ and $N_{\theta}^{(2)}$ use identical U‑Net architectures from (PnPUNet_huang2021variational) with the hyperparameters shown.
GPS Diffusion Architecture
The GPS diffusion module implements the velocity correction term $\mathbf{v}_{\text{diff}}$ using a graph transformer architecture as described in Section 3. We adapted the implementation from the PyTorch Geometric repository [^3] https://github.com/pyg-team/pytorch˙geometric/blob/master/examples/graph˙gps.py. Table 7 presents the hyperparameters used for each dataset in our GPS module experiments. The module:
@@
Table 6: ADM $\mathbf{v}_{\text{react}}$ configurations taken from (dao2023flow). The CelebA-HQ configuration is the one obtained by testing the corresponding checkpoint from (dao2023flow).
Dataset $\mathbf{v}_{\text{react}}$ Hidden Dim Layers PE Dim Walk Length Time Channels
@@
Table 7: Architectural hyperparameters for GPS‐based diffusion term ($\mathbf{v}_{\text{diff}}$). All models used multihead attention with 4 heads and time embeddings with 8 learnable frequencies.
B.4 Training Hyperparameters
We evaluated GFM using two reaction term architectures: DiT-L/2 and ADM U-Net. For both, we initialized $\mathbf{v}_{\text{react}}$ with corresponding architectures and checkpoints from (dao2023flow), then jointly trained the reaction and graph correction (“diffusion term”) components. For AFHQ-Cat experiments, we used the CelebA-HQ checkpoints to initialize the reaction terms. For evaluation of performance metrics, FID scores were monitored every 33K iterations, with training terminating after 8080K iterations without improvement. We saved models with the best FID scores and evaluated them using 5050K generated samples. For CelebA-HQ comparisons, we obtained higher, i.e., worse FID scores than reported in (dao2023flow) when evaluating their provided checkpoints on our machines. To establish fair baselines, we trained these checkpoints for an additional 100100K iterations using the training settings in (dao2023flow) and used the models with the lowest FID scores as the baseline models for the corresponding DiT and ADM experiments.
Tables 8 and 9 detail the training hyperparameters for our GFM variants with DiT-L/2 and ADM U-Net backbones. All models used used the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The original training settings for the (dao2023flow) checkpoints are provided in Tables 10 and 11. Similar to (dao2023flow), a cosine annealing scheduler was used for the learning rate. For ablation studies in Table 3 with identity adjacency matrices, all other settings (besides their adjacency matrices) remained identical to their corresponding graph-based experiments.
Seeds
We set the initial seed during training to $s_{0}=0$. Whenever we generate images to compute FID at epoch $e$, we reset the seed to $s=s_{0}+10^{6}+e$. Finally, for the post-training evaluation on 50,000 generated samples, we fix the seed to $s=42$.
$\mathbf{v}_{\text{diff}}$ Parameter LSUN Church FFHQ LSUN Bedroom AFHQ‑Cat CelebA‑HQ
@@ B.4 Training Hyperparameters
Batch size $42$ $32$ $32$ 2020 $32$
Training iter. 4040K 4040K 6060K 6060K 6060K
# GPUs 11 11 11 11 11
@@
Training iter. 4040K 4040K 4040K 6060K 6060K
# GPUs 11 11 11 11 11
Table 8: Training hyperparameters when $\mathbf{v}_{\text{react}}$ is DiT‑L/2 across datasets. All models used the AdamW optimizer with $\beta_{1}=0.9,\;\beta_{2}=0.999$.
$\mathbf{v}_{\text{diff}}$ Parameter LSUN Church FFHQ LSUN Bedroom AFHQ‑Cat CelebA‑HQ
@@
Training iter. 5050K 120120K 4040K 200200K 6060K
# GPUs 11 11 11 11 11
@@
Training iter. 5050K 120120K 4040K 200200K 6060K
# GPUs 11 11 11 11 11
Table 9: Training hyperparameters when $\mathbf{v}_{\text{react}}$ is ADM (dao2023flow) across datasets. All models used the AdamW optimizer with $\beta_{1}=0.9,\;\beta_{2}=0.999$.
@@
# GPUs 22 11 11
# training days $1.3$ $4.4$ 6.9& 7.36.9\;\&\;7.3
Table 10: Training hyperparameters for the ADM checkpoints used as $\mathbf{v}_{\text{react}}$, reproduced exactly from (dao2023flow).
@@
# GPUs 11 11 2& 12\;\&\;1
# training days $5.8$ $5.5$ 6.1& 12.56.1\;\&\;12.5
Table 11: Training hyperparameters for the DiT checkpoints used as $\mathbf{v}_{\text{react}}$, reproduced exactly from (dao2023flow).
Appendix C Ablation Study: Diffusion-Reaction with KNN
In this ablation, we swap out our attention‐based adjacency for a simple kk‐nearest‐neighbour (KNN) graph in latent space: for each node, we connect to its $k$ closest latents and assign each edge a learnable weight, so that in the diffusion term, information is aggregated strictly from a node’s top‐$k$ neighbors instead of via attention scores. Concretely, we implement the diffusion term as a graph Laplacian multiplied by a per-channel diffusivity vector, mirroring the physical reaction–diffusion system in Equation 6. The overall velocity field remains the sum of the original flow-matching “reaction” network and this KNN-based “diffusion” operator. Let $\mathbf{X}\in\mathbb{R}^{B\times C\times H\times W}$ be the batch of latent tensors at time $t$. This network implements:
$\begin{split}\frac{d\mathbf{X}}{dt}(t)&=\mathbf{v}_{\mathrm{react}}\bigl(\mathbf{X}(t),t\bigr)\\[6.0pt] &\quad+\;\underbrace{\kappa_{\theta}\bigl(\mathbf{X}(t),t\bigr)}_{\begin{subarray}{c}\text{per‐channel}\\ \text{diffusivity}\end{subarray}}\;\odot\;\bigl[\mathbf{L}(\mathbf{X}(t))\,\mathbf{X}_{\mathrm{flat}}(t)\bigr]\,\end{split}$ (9)